OpenStudy (anonymous):

underwater, pressure increases 4.3 pounds per square inch(psi) for every ten feet you decend. This is represented by the equation p=0.43d+14.7, where p is the psi and d is the depth in ft. frind the pressure at 200 feet. a. 215.13psi b.860 psi c.11.7 psi

OpenStudy (anonymous):

for c i ment 100.7 psi

OpenStudy (ash2326):

given p=0.43d+14.7 we need d to find pressure , d is given as 200 p=0.43*200+14.7 p=86+14.7 p=100.7 psi option c